# Network-Optimised Spiking Neural Network for Event-Driven Networking

### v3

@@ Appendix I Stochastic arrivals: shot noise model and calibration
$0\;=\;f_{\mathrm{sat}}(v_{i}^{*})\;+\;\bigl(\beta-\lambda-\chi\bigr)\,v_{i}^{*}\;+\;\gamma\;+\;\chi v_{\mathrm{rest}}\;-\;u_{i}^{*}\;+\;I_{i}^{*},$ (50)
@@
$f_{\mathrm{sat}}(v_{i}^{*})\;+\;\Bigl(\beta-\lambda-\chi-\frac{ab}{a+\mu}\Bigr)\,v_{i}^{*}\;+\;\gamma\;+\;\chi v_{\mathrm{rest}}\;+\;I_{i}^{*}\;=\;0.$ (51)
@@
If $D_{i}\geq 0$ and the admissible root lies in $[0,1]$, it approximates $v_{i}^{*}$ and clarifies how the operating point moves with $C$. In networks, service and damping must outpace the steepest small-signal gain of $f_{\mathrm{sat}}$; the corollary in M provided a direct algebraic check. In the quadratic approximation, the sign of $D_{i}$ in (54) separates three regimes: $D_{i}<0$ (no real root, parameter mismatch), $D_{i}=0$ (tangent balance at a single $v_{i}^{*}$), and $D_{i}>0$ (two mathematical roots, of which the physically admissible one lies in the queue range and satisfies the stability test below).
@@
The monotonicity criterion introduced earlier applies with $L$ as in (52): a sufficiently positive net slope $f^{\prime}_{\mathrm{sat}}(v)+L$ over the admissible interval (for example $v\in[0,1]$) ensures a unique solution. For the saturating choice
@@
$\inf_{v\in[0,1]}\bigl(f^{\prime}_{\mathrm{sat}}(v)+L\bigr)\;>\;0,$ (57)
then $F$ is strictly increasing on $[0,1]$ and has at most one root there. If in addition $F(0)\,F(1)\leq 0$, a unique $v_{i}^{*}\in[0,1]$ exists. Condition (57) is met whenever
$L\;>\;-\,\frac{3\sqrt{3}}{8}\,\frac{\alpha}{\sqrt{\kappa}},$ (58)
which is the small-signal service-dominance requirement: linear damping $\lambda+\chi+\tfrac{ab}{a+\mu}$ must exceed the maximal small-signal excitability slope.
@@
Equation (51) balances offered load against service and damping. The term $I_{i}^{*}$ aggregates graph-local arrivals and policy weights; $L$ collects all linear drains. The sensitivity to slowly varying load follows by implicit differentiation:
@@
Hence the DC gain is stable and interpretable. It shrinks as saturation increases $(\kappa\uparrow)$ or as service and damping increase $(\lambda,\chi,ab/(a+\mu)\uparrow)$, which matches the operational goal of keeping the queue operating point insensitive to modest load drift.
K.2 Jacobian and linear stability (details)
@@
Lemma 1 (Monotone stabilisation by saturation).
Let $f_{\mathrm{sat}}(v)=\dfrac{\alpha v^{2}}{1+\kappa v^{2}}$ with $\alpha>0$, $\kappa>0$, and let $v^{*}\geq 0$ be a subthreshold equilibrium. Then
$\frac{\partial}{\partial\kappa}\,f^{\prime}(v^{*})\;=\;-\,\frac{4\alpha(v^{*})^{3}}{\bigl(1+\kappa(v^{*})^{2}\bigr)^{3}}\;\leq\;0,$ (67)
with equality only at $v^{*}=0$. Consequently,
$\begin{split}\frac{\partial T_{i}}{\partial\kappa}\;=\;\frac{\partial f^{\prime}(v^{*})}{\partial\kappa}\;<\;0\quad(v^{*}>0),\\ \frac{\partial\Delta_{i}}{\partial\kappa}\;=\;-\,(a+\mu)\,\frac{\partial f^{\prime}(v^{*})}{\partial\kappa}\;>\;0\quad(v^{*}>0),\end{split}$ (68)
so increasing $\kappa$ strictly decreases the trace and strictly increases the determinant, enlarging the region $\{T_{i}<0,\ \Delta_{i}>0\}$.
@@
Differentiate $f^{\prime}(v)=\dfrac{2\alpha v}{(1+\kappa v^{2})^{2}}$ with respect to $\kappa$:
$\frac{\partial}{\partial\kappa}f^{\prime}(v)\;=\;-\,\frac{4\alpha v^{3}}{\bigl(1+\kappa v^{2}\bigr)^{3}},$ (69)
which is nonpositive for $v\geq 0$ and strictly negative for $v>0$. Substituting into (65)–(66) gives (68). ∎
Corollary 1.1 (Network threshold increases with saturation).
@@
$\bar{d}\;:=\;f^{\prime}(v^{*})+\beta-\lambda-\chi,\qquad k^{\star}\;=\;\min\Bigl\{(a+\mu)-\bar{d},\ \frac{ab}{a+\mu}-\bar{d}\Bigr\},\qquad g_{\star}\;\approx\;\frac{k^{\star}}{\rho(W)}.$ (70)
Since $\dfrac{\partial\bar{d}}{\partial\kappa}=\dfrac{\partial f^{\prime}(v^{*})}{\partial\kappa}<0$ for $v^{*}>0$, one has $\dfrac{\partial k^{\star}}{\partial\kappa}>0$ and therefore $\dfrac{\partial g_{\star}}{\partial\kappa}>0$ on the active branch of $k^{\star}$. Increasing $\kappa$ raises the coupling needed to destabilise the network, which matches the bounded-excitability design goal.
In networking terms, making the excitability saturate earlier reduces small-signal gain and raises damping, which protects the node against oscillations when operated near capacity.
@@
Let $\lambda_{1,2}$ be the eigenvalues of the local Jacobian $J_{i}$ defined in K.2. When the Routh–Hurwitz conditions in (63) hold and $T_{i}^{2}>4\Delta_{i}$, the node is an overdamped node with two real negative modes. When $T_{i}^{2}<4\Delta_{i}$, it is underdamped with complex conjugate modes whose common decay rate is $\tfrac{1}{2}T_{i}$. The dominant linear return time is
@@
K.3 Network coupling and global stability (details)
@@
$I_{i}^{*}\;=\;\sum_{j}w_{ij}\,S_{j}^{*},\qquad\max_{i}I_{i}^{*}\;\leq\;\|W\|_{\infty}\,\|S^{*}\|_{\infty},\qquad\|I_{i}^{*}\|_{2}\leq\;\|W\|_{2}\,\|S^{*}\|_{2}\ \$(73)
@@
$C\;\leq\;\frac{\alpha}{\kappa}\quad\Longrightarrow\quad\text{a nonnegative equilibrium exists when }L\leq 0.$ (76)
@@
Using the legacy “starter” values (kept here for continuity) $\alpha=0.02$, $\beta=0.5$, $\gamma=0.05$, $\lambda=0.2$, $a=0.05$, $b=0.5$, $\mu=0.01$, $\chi=0.05$, and $I_{\max}=0.10$,
$\frac{ab}{a+\mu}=\frac{0.025}{0.06}\approx 0.4167,\quad L=0.5-0.2-0.05-0.4167\approx-0.1667,$
$\frac{L^{2}}{4\alpha}\approx\frac{0.02778}{0.08}=0.34722,\quad\Delta_{\mathrm{op}}\approx 0.34722-(0.05+0.10)=0.19722>0.$
At $I_{\max}=0.30$ the margin is $\Delta_{\mathrm{op}}\approx 0$, and raising $\chi$ from $0.05$ to $0.30$ changes $L$ from 0.1667-0.1667 to 0.4167-0.4167, yielding $\Delta_{\mathrm{op}}\approx 1.82$—a substantial safety increase. Note: these numbers are purely illustrative for the quadratic surrogate; in the main experiments we use the admissible ranges in Table 4. The linear dependence on $I_{\max}$ and the monotone effect of $\chi$ hold in either case.
@@
K.4 Block Jacobian, network spectrum, and stability threshold (details)
@@
$\mathcal{J}\;=\;\begin{bmatrix}D+G&-\mathbf{I}_{N}\\ aB&-(a+\mu)\mathbf{I}_{N}\end{bmatrix},\qquad D=\operatorname{diag}\!\Bigl(f^{\prime}(v_{i}^{*})+\beta-\lambda-\chi\Bigr),\quad B=b\mathbf{I}_{N}.$ (81)
@@
$J_{\mathrm{eff}}(k)\;=\;\begin{bmatrix}\bar{d}+k&-1\\ ab&-(a+\mu)\end{bmatrix},\qquad k\;:=\;g\,\rho(W),\qquad\bar{d}\;:=\;\frac{1}{\|\phi\|_{2}^{2}}\sum_{i}\phi_{i}^{2}\bigl[f^{\prime}(v_{i}^{*})+\beta-\lambda-\chi\bigr].$ (82)
@@
Using $\Lambda:=\lambda+\chi+\frac{ab}{a+\mu}-\beta$ (cf. (99)), the threshold condition can be written as
$g\,\rho(W)\,f^{\prime}_{\mathrm{sat}}(v^{*})\;<\;\Lambda,$ (84)
which matches the stability proxy introduced in the scaling discussion. While the Perron-mode reduction yields tight thresholds in practice, conservative certificates can be obtained via Gershgorin discs and norm bounds; see methods K.5 for the resulting inequalities and their comparison to measured $g_{\star}$ in Table 6.
@@
The rescaled thresholds are not perfectly flat. They sit slightly above $k^{\star}$ across $\rho(W)$, with larger deviations in smaller or more heterogeneous graphs. Two mechanisms explain this percent-level bias. First, nearby non-Perron modes and row–sum variability shift the leading root when the Perron vector localises on hubs, a common feature of heavy-tailed topologies. Second, discrete sweeps of $\Re\lambda_{\max}(g)$ and mild non-normal amplification nudge the numerical crossing. These are finite-size corrections, not a change of scaling, and they diminish with larger $N$ or more homogeneous weights. For conservative guarantees during early deployment, replace $\rho(W)$ by norm bounds or apply Gershgorin discs to $\mathcal{J}$ to obtain stricter, topology-aware thresholds. For deployments requiring analytic guarantees, methods K.5 derives Gershgorin and $\|W\|_{\infty}$-based bounds and reports their conservatism relative to measured $g^{\star}$ (Table 6). A detailed discussion on the bifurcation structure for NOS, and its operational interpretation, is given in L.
K.5 Conservative stability via Gershgorin and norm bounds
Let $G=gW$ with $W_{ii}=0$ and $\rho(W)$ the spectral radius after normalisation (we use $\rho(W)=1$ in experiments). Linearising (2)–(3) at $(v^{*},u^{*})$ gives a $2N\times 2N$ block Jacobian $\mathcal{J}$. For the top block rows 1:N1{:}N,
$\bar{d}_{i}:=f^{\prime}_{\mathrm{sat}}(v_{i}^{*})+\beta-\lambda-\chi,$
is the diagonal entry, and the off-diagonals are $gW_{ij}$ for $j\neq i$, plus the coupling to $u_{i}$ with derivative 1-1. Thus the Gershgorin centre and radius are
$c_{i}=\bar{d}_{i},\qquad R_{i}=1+|g|\sum_{j\neq i}|W_{ij}|.$
A sufficient condition for those discs to lie strictly in the left half-plane is
$\Re(c_{i})+R_{i}<0\quad\text{for all }i,$
which yields
$|g|\;<\;\min_{i}\frac{-\Re(\bar{d}_{i})-1}{\sum_{j\neq i}|W_{ij}|}.$
For the bottom block rows N+1:2NN{+}1{:}2N, the centre is $(a+\mu)$ and the radius is $|ab|$, so a sufficient condition is
$(a+\mu)+|ab|<0.$
Using $\sum_{j\neq i}|W_{ij}|\leq\|W\|_{\infty}$ gives the norm variant
$|g|\;<\;\frac{-\min_{i}\Re(\bar{d}_{i})-1}{\|W\|_{\infty}},\qquad\text{and}\qquad|ab|<a+\mu.$
These conditions are sufficient (not necessary) and are conservative relative to observed thresholds.
@@
In finite graphs the transition is slightly rounded but sharpens with size. The measured coupling threshold aligns with the Perron-mode prediction $k^{\star}$ (cf. (83)), with small upward deviations when the Perron vector localises on hubs or when secondary modes lie close in the spectrum. These are percent-level corrections that diminish in larger or more homogeneous topologies. Practically, one can monitor the rescaled margin $\rho(W)\,g-k^{\star}$: values near zero predict imminent entrainment of queues and the appearance of coherent oscillations along the dominant influence path. Online-Methods O quantifies network coherence via an order parameter $R(t)$, confirming that the onset clusters near $k^{\star}$ and sharpens with $N$ (Fig. 15, Table 8).

### v4

@@ Appendix I Stochastic arrivals: shot noise model and calibration
$\begin{split}0\;=\;f_{\mathrm{sat}}(v_{i}^{*})\;+\;\bigl(\beta-\lambda-\chi\bigr)\,v_{i}^{*}\;+\;\gamma\\ \;+\;\chi v_{\mathrm{rest}}\;-\;u_{i}^{*}\;+\;I_{i}^{*},\end{split}$ (49)
@@
$\begin{split}f_{\mathrm{sat}}(v_{i}^{*})\;+\;\Bigl(\beta-\lambda-\chi-\frac{ab}{a+\mu}\Bigr)\,v_{i}^{*}\;+\;\gamma\\ \;+\chi v_{\mathrm{rest}}\;+\;I_{i}^{*}\;=\;0.\end{split}$ (50)
@@
If $D_{i}\geq 0$ and the admissible root lies in $[0,1]$, it approximates $v_{i}^{*}$ and clarifies how the operating point moves with $C$. In networks, a convenient sufficient condition for a unique operating point is that the net small-signal slope is negative on $[0,1]$, i.e. $f^{\prime}_{\mathrm{sat}}(v)+L<0$ throughout the admissible range; the corollary in M provides a direct algebraic check. In the quadratic approximation, the sign of $D_{i}$ in (53) separates three regimes: $D_{i}<0$ (no real root under the quadratic approximation), $D_{i}=0$ (tangent balance at a single $v_{i}^{*}$), and $D_{i}>0$ (two mathematical roots, of which the physically admissible one lies in the queue range and is then checked against the monotonicity and stability conditions below).
@@
The monotonicity criterion introduced earlier applies with $L$ as in (51): a sufficiently negative net slope $f^{\prime}_{\mathrm{sat}}(v)+L$ over the admissible interval (for example $v\in[0,1]$) ensures a unique solution. For the saturating choice
@@
$\sup_{v\in[0,1]}\bigl(f^{\prime}_{\mathrm{sat}}(v)+L\bigr)\;<\;0,$ (56)
then $F$ is strictly decreasing on $[0,1]$ and has at most one root there. If in addition $F(0)\,F(1)\leq 0$, a unique $v_{i}^{*}\in[0,1]$ exists. Condition (56) is met whenever
$L\;<\;-\,\frac{3\sqrt{3}}{8}\,\frac{\alpha}{\sqrt{\kappa}},$ (57)
which is the small-signal service-dominance requirement: linear damping $\lambda+\chi+\tfrac{ab}{a+\mu}$ must dominate the maximal small-signal excitability slope (offsetting any positive linear gain $\beta$).
@@
Equation (50) balances offered load against service and damping. The term $I_{i}^{*}$ aggregates graph-local arrivals and policy weights; $L$ collects all linear contributions. The sensitivity to slowly varying load follows by implicit differentiation:
@@
Hence the DC gain is well-defined. Under the monotonicity condition (56), the denominator is strictly negative on $[0,1]$, so $\partial v_{i}^{*}/\partial I_{i}^{*}>0$ as expected: increasing offered load increases the operating queue level. The gain shrinks as saturation strengthens $(\kappa\uparrow)$ or as service and damping increase $(\lambda,\chi,ab/(a+\mu)\uparrow)$, matching the operational goal of keeping the operating point insensitive to modest load drift.
K.2 Derivation note: $\delta I\approx G\,\delta v$
Let the coupling current be $I(v)=g\,W\,f_{\mathrm{sat}}(v)$ (applied with the stated delay). Linearising about $v^{*}$ gives
$\delta I\;=\;g\,W\,f^{\prime}_{\mathrm{sat}}(v^{*})\,\delta v\;+\;\mathcal{O}(\|\delta v\|^{2}).$ (59)
Taking an induced norm yields the scalar bound
$\|\delta I\|\;\leq\;g\,\|W\|\,|f^{\prime}_{\mathrm{sat}}(v^{*})|\,\|\delta v\|\;=:\;G\,\|\delta v\|,$ (60)
so we write $\delta I\approx G\,\delta v$ as shorthand, with
$G\;:=\;g\,\|W\|\,|f^{\prime}_{\mathrm{sat}}(v^{*})|.$ (61)
K.3 Jacobian and linear stability (details)
@@
From local Jacobian to a network-level proxy.
To connect (65) to the graph-level proxy used in the main text, consider homogeneous coupling where small drive perturbations satisfy $\delta I\approx G\,\delta v$ with $G=gW$ and Perron eigenvalue $\rho(W)$. Define
$\bar{d}\;:=\;f^{\prime}(v^{*})+\beta-\lambda-\chi.$ (66)
Stacking states $(v,u)\in\mathbb{R}^{2N}$ gives the block Jacobian
$\mathcal{J}\;=\;\begin{bmatrix}\bar{d}\,\mathbf{I}_{N}+gW&-\mathbf{I}_{N}\\ ab\,\mathbf{I}_{N}&-(a+\mu)\mathbf{I}_{N}\end{bmatrix}$. (67)
Along any eigenmode of $W$ with eigenvalue $\omega$, the dynamics reduce to an effective $2\times 2$ Jacobian
$J_{\mathrm{eff}}(\omega)\;=\;\begin{bmatrix}\bar{d}+g\omega&-1\\ ab&-(a+\mu)\end{bmatrix}$. (68)
The least stable mode uses $\omega=\rho(W)$, so write $k:=g\rho(W)$. Routh–Hurwitz on (68) yields the two conditions
$k\;<\;(a+\mu)-\bar{d},\qquad k\;<\;\frac{ab}{a+\mu}-\bar{d}.$ (69)
Now define the net drain
$\Lambda\;:=\;\lambda+\chi+\frac{ab}{a+\mu}-\beta.$ (70)
Using (66), the determinant branch in (69) is exactly
$k+f^{\prime}(v^{*})\;<\;\Lambda,\qquad\text{i.e.}\qquad g\,\rho(W)\;<\;\Lambda-f^{\prime}(v^{*}),$ (71)
which is the additive form used for the network stability proxy.
Remark (normalised reporting).
Any ratio form (for example $(k+f^{\prime}(v^{*}))/\Lambda$) is only a normalised headroom score. The stability condition is the additive inequality (71).
@@
Lemma 1 (Monotone stabilisation by saturation (fixed operating point)).
Let $f_{\mathrm{sat}}(v)=\dfrac{\alpha v^{2}}{1+\kappa v^{2}}$ with $\alpha>0$ and $\kappa>0$. For any fixed $v\geq 0$,
$\frac{\partial}{\partial\kappa}\,f^{\prime}_{\mathrm{sat}}(v)\;=\;-\,\frac{4\alpha v^{3}}{\bigl(1+\kappa v^{2}\bigr)^{3}}\;\leq\;0,$ (75)
with equality only at $v=0$. Consequently, for the Jacobian (62) evaluated at a fixed operating point $v^{*}>0$, increasing $\kappa$ strictly decreases the trace $T_{i}$ and strictly increases the determinant $\Delta_{i}$, thereby enlarging the region $\{T_{i}<0,\ \Delta_{i}>0\}$ when re-evaluated at that same $v^{*}$.
@@
Differentiate $f^{\prime}_{\mathrm{sat}}(v)=\dfrac{2\alpha v}{(1+\kappa v^{2})^{2}}$ with respect to $\kappa$ to obtain (75). For fixed $v^{*}$, the $\kappa$-dependence of (62) enters only through $f^{\prime}_{\mathrm{sat}}(v^{*})$, hence
$\begin{split}\frac{\partial T_{i}}{\partial\kappa}\;=\;\frac{\partial f^{\prime}_{\mathrm{sat}}(v^{*})}{\partial\kappa}\;<\;0\quad(v^{*}>0),\\ \frac{\partial\Delta_{i}}{\partial\kappa}\;=\;-(a+\mu)\frac{\partial f^{\prime}_{\mathrm{sat}}(v^{*})}{\partial\kappa}\;>\;0,\end{split}$ (76)
since $a+\mu>0$ and (75) is strict for $v^{*}>0$. ∎
Remark (equilibrium shift with $\kappa$).
If the operating point is taken as the equilibrium branch $v^{*}=v^{*}(\kappa)$ rather than held fixed, then the relevant quantities use total derivatives. In particular,
$\frac{d}{d\kappa}\,f^{\prime}_{\mathrm{sat}}(v^{*}(\kappa),\kappa)\;=\;\frac{\partial f^{\prime}_{\mathrm{sat}}}{\partial\kappa}(v^{*},\kappa)\;+\;f^{\prime\prime}_{\mathrm{sat}}(v^{*},\kappa)\,\frac{dv^{*}}{d\kappa},$ (77)
so $\frac{dT_{i}}{d\kappa}=\frac{d}{d\kappa}f^{\prime}_{\mathrm{sat}}(v^{*})$ and $\frac{d\Delta_{i}}{d\kappa}=-(a+\mu)\frac{d}{d\kappa}f^{\prime}_{\mathrm{sat}}(v^{*})$.
Moreover, $dv^{*}/d\kappa$ follows from the equilibrium relation (Eq. (58)). Writing $F(v,\kappa):=f_{\mathrm{sat}}(v,\kappa)+Lv+C=0$ with $L:=\beta-\lambda-\chi-\frac{ab}{a+\mu}$, the implicit-function theorem gives
$\frac{dv^{*}}{d\kappa}\;=\;-\,\frac{\partial_{\kappa}F(v^{*},\kappa)}{\partial_{v}F(v^{*},\kappa)}\;=\;-\,\frac{\partial_{\kappa}f_{\mathrm{sat}}(v^{*},\kappa)}{f^{\prime}_{\mathrm{sat}}(v^{*},\kappa)+L},$ (78)
where $\partial_{\kappa}f_{\mathrm{sat}}(v,\kappa)=-\dfrac{\alpha v^{4}}{(1+\kappa v^{2})^{2}}<0$ for $v>0$. Thus the pointwise monotonicity (75) holds unconditionally at fixed $v^{*}$, while monotonic behaviour along $v^{*}(\kappa)$ depends on the additional term in (77). In the paper we use (75) as an engineering fact: for a given observed subthreshold $v^{*}$, larger $\kappa$ reduces the small-signal slope and widens the Routh–Hurwitz margins when re-evaluated at that operating point.
Corollary 1.1 (Network threshold increases with saturation (fixed operating point)).
@@
$\begin{split}\bar{d}\;:=\;f^{\prime}_{\mathrm{sat}}(v^{*})+\beta-\lambda-\chi,\\ k^{\star}\;=\;\min\Bigl\{(a+\mu)-\bar{d},\ \frac{ab}{a+\mu}-\bar{d}\Bigr\},\qquad g_{\star}\;\approx\;\frac{k^{\star}}{\rho(W)}.\end{split}$ (79)
Equivalently, with $\Lambda$ from (70), the determinant branch is $k<\Lambda-f^{\prime}_{\mathrm{sat}}(v^{*})$, matching (71). Holding $v^{*}$ fixed, (75) implies $\partial\bar{d}/\partial\kappa<0$, hence $\partial k^{\star}/\partial\kappa>0$ on the active branch, and therefore $\partial g_{\star}/\partial\kappa>0$.
@@
Let $\lambda_{1,2}$ be the eigenvalues of the local Jacobian $J_{i}$ defined in K.3. When the Routh–Hurwitz conditions in (65) hold and $T_{i}^{2}>4\Delta_{i}$, the node is an overdamped node with two real negative modes. When $T_{i}^{2}<4\Delta_{i}$, it is underdamped with complex conjugate modes whose common decay rate is $\tfrac{1}{2}T_{i}$. The dominant linear return time is
@@
K.4 Delay-aware Perron-mode stability (DDE extension)
For homogeneous coupling with a representative delay $\tau$ on the dominant mode, the linearised Perron-mode dynamics become a $2\times 2$ delay system:
$\dot{\delta v}(t)$ $=\bar{d}\,\delta v(t)-\delta u(t)+k\,\delta v(t-\tau),$ (82)
$\dot{\delta u}(t)$ $=ab\,\delta v(t)-(a+\mu)\,\delta u(t),$ (83)
where $\bar{d}=f^{\prime}(v^{*})+\beta-\lambda-\chi$ and $k=g\rho(W)$. The characteristic equation is
$\bigl(s-\bar{d}-ke^{-s\tau}\bigr)\bigl(s+(a+\mu)\bigr)+ab=0.$ (84)
Setting $s=i\Omega$ gives the boundary of stability in the standard frequency-domain form. Define
$N(\Omega)=\bigl(i\Omega+(a+\mu)\bigr)\bigl(i\Omega-\bar{d}\bigr)+ab.$
Then (84) implies
$k\,e^{-i\Omega\tau}=\frac{N(\Omega)}{i\Omega+(a+\mu)}.$ (85)
Hence the critical coupling at frequency $\Omega$ is
$k(\Omega)=\frac{|N(\Omega)|}{\sqrt{\Omega^{2}+(a+\mu)^{2}}},$ (86)
and admissible delays satisfy the phase relation
$\tau=\frac{1}{\Omega}\Bigl(\arg(i\Omega+(a+\mu))-\arg(N(\Omega))+2\pi m\Bigr),\quad m\in\mathbb{Z}.$ (87)
For a given $\tau$, one finds $\Omega$ satisfying (87) (for some integer $m$) and evaluates $k_{\mathrm{crit}}(\tau)=\min k(\Omega)$ over the feasible crossings. When $\tau=0$, the $\Omega\to 0$ limit of (86) recovers the undelayed determinant branch $k<\frac{ab}{a+\mu}-\bar{d}$, which is equivalent to $k+f^{\prime}(v^{*})<\Lambda$ with $\Lambda=\lambda+\chi+\frac{ab}{a+\mu}-\beta$. Empirically, increasing $\tau$ shifts $k_{\mathrm{crit}}(\tau)$ downward, matching the boundary shift observed in Fig. 3(a).
K.5 Network coupling and global stability (details)
@@
$\begin{split}I_{i}^{*}\;=\;\sum_{j}w_{ij}\,S_{j}^{*},\\ \max_{i}I_{i}^{*}\;\leq\;\|W\|_{\infty}\,\|S^{*}\|_{\infty},\\ \|I_{i}^{*}\|_{2}\leq\;\|W\|_{2}\,\|S^{*}\|_{2}\ \ \end{split}$ (88)
@@
$C\;\leq\;\frac{\alpha}{\kappa}\;\rightarrow\;\text{a nonnegative equilibrium exists when }L\leq 0.$ (91)
@@
Using the legacy “starter” values (kept here for continuity) $\alpha=0.02$, $\beta=0.5$, $\gamma=0.05$, $\lambda=0.2$, $a=0.05$, $b=0.5$, $\mu=0.01$, $\chi=0.05$, and $I_{\max}=0.10$, $\frac{ab}{a+\mu}=\frac{0.025}{0.06}\approx 0.4167,$ $L=0.5-0.2-0.05-0.4167\approx-0.1667,$ $\frac{L^{2}}{4\alpha}\approx\frac{0.02778}{0.08}=0.34722,\quad\Delta_{\mathrm{op}}\approx 0.34722-(0.05+0.10)=0.19722>0.$ At $I_{\max}=0.30$ the margin is $\Delta_{\mathrm{op}}\approx 0$, and raising $\chi$ from $0.05$ to $0.30$ changes $L$ from 0.1667-0.1667 to 0.4167-0.4167, yielding $\Delta_{\mathrm{op}}\approx 1.82$—a substantial safety increase. Note: these numbers are purely illustrative for the quadratic surrogate; in the main experiments we use the admissible ranges in Table 4. The linear dependence on $I_{\max}$ and the monotone effect of $\chi$ hold in either case.
@@
K.6 Block Jacobian, network spectrum, and stability threshold (details)
@@
$\begin{split}\mathcal{J}\;=\;\begin{bmatrix}D+G&-\mathbf{I}_{N}\\ aB&-(a+\mu)\mathbf{I}_{N}\end{bmatrix},\\ D=\operatorname{diag}\!\Bigl(f^{\prime}(v_{i}^{*})+\beta-\lambda-\chi\Bigr),\quad B=b\mathbf{I}_{N}.\end{split}$ (96)
@@
$\begin{split}J_{\mathrm{eff}}(k)\;=\;\begin{bmatrix}\bar{d}+k&-1\\ ab&-(a+\mu)\end{bmatrix},\\ k\;:=\;g\,\rho(W),\\ \bar{d}\;:=\;\frac{1}{\|\phi\|_{2}^{2}}\sum_{i}\phi_{i}^{2}\bigl[f^{\prime}(v_{i}^{*})+\beta-\lambda-\chi\bigr].\end{split}$ (97)
@@
Using $\Lambda:=\lambda+\chi+\frac{ab}{a+\mu}-\beta$ (cf. (114)) and $\bar{d}\approx f^{\prime}_{\mathrm{sat}}(v^{*})+\beta-\lambda-\chi$, the determinant branch of (98) gives
$g\,\rho(W)\;<\;\Lambda-f^{\prime}_{\mathrm{sat}}(v^{*}),$ (99)
equivalently $g\,\rho(W)+f^{\prime}_{\mathrm{sat}}(v^{*})<\Lambda$, which matches the stability proxy introduced in the scaling discussion. While the Perron-mode reduction yields tight thresholds in practice, conservative certificates can be obtained via Gershgorin discs and norm bounds; see methods K.7 for the resulting inequalities and their comparison to measured $g_{\star}$ in Table 6.
@@
The rescaled thresholds are not perfectly flat. They sit slightly above $k^{\star}$ across $\rho(W)$, with larger deviations in smaller or more heterogeneous graphs. Two mechanisms explain this percent-level bias. First, nearby non-Perron modes and row–sum variability shift the leading root when the Perron vector localises on hubs, a common feature of heavy-tailed topologies. Second, discrete sweeps of $\Re\lambda_{\max}(g)$ and mild non-normal amplification nudge the numerical crossing. These are finite-size corrections, not a change of scaling, and they diminish with larger $N$ or more homogeneous weights. For conservative guarantees during early deployment, replace $\rho(W)$ by norm bounds or apply Gershgorin discs to $\mathcal{J}$ to obtain stricter, topology-aware thresholds. For deployments requiring analytic guarantees, methods K.7 derives Gershgorin and $\|W\|_{\infty}$-based bounds and reports their conservatism relative to measured $g^{\star}$ (Table 6). A detailed discussion on the bifurcation structure for NOS, and its operational interpretation, is given in L.
K.7 Conservative stability via Gershgorin and norm bounds
Let $G=gW$ with $W_{ii}=0$ and $\rho(W)$ the spectral radius after normalisation (we use $\rho(W)=1$ in experiments). Linearising (2)–(3) at $(v^{*},u^{*})$ gives a $2N\times 2N$ block Jacobian $\mathcal{J}$. For the top block rows 1:N1{:}N, $\bar{d}_{i}:=f^{\prime}_{\mathrm{sat}}(v_{i}^{*})+\beta-\lambda-\chi,$ is the diagonal entry, and the off-diagonals are $gW_{ij}$ for $j\neq i$, plus the coupling to $u_{i}$ with derivative 1-1. Thus the Gershgorin centre and radius are $c_{i}=\bar{d}_{i},\qquad R_{i}=1+|g|\sum_{j\neq i}|W_{ij}|.$ A sufficient condition for those discs to lie strictly in the left half-plane is $\Re(c_{i})+R_{i}<0\quad\text{for all }i,$ which yields $|g|\;<\;\min_{i}\frac{-\Re(\bar{d}_{i})-1}{\sum_{j\neq i}|W_{ij}|}.$ For the bottom block rows N+1:2NN{+}1{:}2N, the centre is $(a+\mu)$ and the radius is $|ab|$, so a sufficient condition is $(a+\mu)+|ab|<0.$ Using $\sum_{j\neq i}|W_{ij}|\leq\|W\|_{\infty}$ gives the norm variant $|g|\;<\;\frac{-\min_{i}\Re(\bar{d}_{i})-1}{\|W\|_{\infty}},\qquad\text{and}\qquad|ab|<a+\mu.$ These conditions are sufficient (not necessary) and are conservative relative to observed thresholds.
@@
In finite graphs the transition is slightly rounded but sharpens with size. The measured coupling threshold aligns with the Perron-mode prediction $k^{\star}$ (cf. (98)), with small upward deviations when the Perron vector localises on hubs or when secondary modes lie close in the spectrum. These are percent-level corrections that diminish in larger or more homogeneous topologies. Practically, one can monitor the rescaled margin $\rho(W)\,g-k^{\star}$: values near zero predict imminent entrainment of queues and the appearance of coherent oscillations along the dominant influence path. The methods O quantifies network coherence via an order parameter $R(t)$, confirming that the onset clusters near $k^{\star}$ and sharpens with $N$ (Fig. 15, Table 8).